Suppose that two objects attract each other with force of 16 units. If the distance between the two objects is doubled, what is the new force of attraction between the two objects


1
Expert's answer
2022-01-16T13:10:00-0500

"F\sim\frac 1{r^2},"

"F_2=\frac{F_1}{2^2}=4."